OpenStudy (anonymous):

Solve the rational equation. Express numbers as integers or simplified fractions. 4/3y^2-12 -6/y+2+1/y-2=0

OpenStudy (***[isuru]***):

i guess it is something like this ' cause its not clear in the way u have type it .. \[\frac{ 4 }{ 3y^{2} - 12 } - \frac{ 6 }{ y + 2} + \frac{ 1 }{ y - 2 }\] now u can write \[\frac{ 4 }{ 3y^{2} - 12 } = \frac{ 4 }{ 3( y^{2} - 4) }\] now using the difference of 2 squares \[\frac{ 4 }{ 3( y^{2} - 4) } = \frac{ 4 }{ 3( y - 2) ( y+ 2) }\] so... u can write ur question in this way now \[\frac{ 4 }{ 3( y- 2)(2 + 2) } - \frac{ 6 }{ y + 2 } + \frac{ 1 }{ y - 2 }\] now take a common denominator... i h think u know how to find it... here it's \[3(y - 2)( y +2)\] now write the whole question as a single fraction \[\frac{ 4 - 6[3(y-2)] + 1[3(y +2) }{ 3(y+2)(y-2) }\] which is \[\frac{ 4 + 18(y-2) + 3(y+2) }{ 3(y-2)(y+2) } = 0\] now u can move the denominator to the other side. since the other side is 0 the whole term will become zero , giving u \[4 + 18(y - 2) + 3(y +2) = 0\] now .. solve this for "y" .. i think u can do that :)

OpenStudy (anonymous):

Would it be 4+18y-36+3y+6=0?

OpenStudy (***[isuru]***):

yep :)

OpenStudy (anonymous):

Would it be -26=21y?

OpenStudy (***[isuru]***):

yep :)